### v1

# Iñigo Alonso & Mirella Lapata School of Informatics

Abstract
While table understanding increasingly relies on pixel-only settings where tables are processed as visual representations, current benchmarks predominantly use synthetic renderings that lack the complexity and visual diversity of real-world tables. Additionally, existing visual table understanding (VTU) datasets offer fixed examples with single visualizations and pre-defined instructions, providing no access to underlying serialized data for reformulation. We introduce Tablet, a large-scale VTU dataset with 4 million examples across 20 tasks, grounded in 2 million unique tables where 88% preserve original visualizations. Each example includes paired image-HTML representations, comprehensive metadata, and provenance information linking back to the source datasets. Fine-tuning vision-language models like Qwen2.5-VL-7B on Tablet improves performance on seen and unseen VTU tasks while increasing robustness on real-world table visualizations. By preserving original visualizations and maintaining example traceability in a unified large-scale collection, Tablet establishes a foundation for robust training and extensible evaluation of future VTU models. [^1] Dataset, code, and other resources will be publicly available.
1 Introduction
The field of table understanding focuses on techniques for representing and interpreting tabular data to support a wide range of practical tasks such as question answering, summarization, and information extraction. Research in this area has traditionally representated tables as structured text, encoding their content and layout through linearized or graph-based representations (see Figure 1b; Herzig et al. 2020; Zhang et al. 2020; Liu et al. 2022). While this unimodal view remains effective in certain domains, many tables found in documents and webpages contain irregular structures, rely on visual formatting (e.g., merged cells, background colors, font variations), or embed multimodal elements such as images (see Figure 1a). Advances in Vision-Language Models (VLMs; Radford et al. 2021; Liu et al. 2023) have provided impetus for treating tables as images, eschewing the step of rendering them as text sequences (like Markdown or HTML). The conceptual simplicity of this approach, coupled with improved performance on several tabular tasks (Alonso et al., 2024; Zhou et al., 2025) has driven significant research interest (Zheng et al., 2024b; Su et al., 2024; Jiang et al., 2025) in Visual Table Understanding (also known as Multimodal Table Understanding). Visual representations of tables are not only merely convenient but in many cases necessary, particularly for VLM agents that interact with the world exclusively through pixels (e.g., on a screen) and must interpret tables directly in their visual form (Deng et al., 2023; Zheng et al., 2024a; Lu et al., 2024).
Despite the growing relevance of VTU, there are few resources that support training models directly on image-based representations of tables. Existing benchmarks like MMTab (Zheng et al., 2024b) consist of web tables (e.g., from Wikipedia which is a common source for many tabular datasets), that are serialized and subsequently rendered as synthetic images (see Figures 1b,c). These images do not preserve the original visual characteristics of the tables and do not reflect the diversity and complexity of real-world layouts. As a result, models trained on such data face a train-test mismatch, since the visual patterns learned from serialized renderings do not generalize well to naturally occurring tables failing to capture critical visual cues like subtle ruling lines, intricate merged cell layouts, background colors, font variations, or embedded images that are inherent to real-world table comprehension (compare Figure 1a and 1c). An exception is WikiDT Shi et al. (2024), which provides access to original table visualizations from web pages but focuses exclusively on table question answering. Likewise, datasets designed for processing PDF documents or screenshots (Zhong et al., 2020; Lu et al., 2023) preserve original table visualizations but support only a single task.
In this work, we introduce Tablet [^2] Our dataset is called Tablet partly as a nod to the Scottish sugary confection which is often cut into uniform squares or rectangles, resembling a table., a large-scale dataset designed to enable vision-language models to learn generalizable skills for table understanding by leveraging visual features from table images. Unlike existing datasets that focus on a single task or use synthetic renderings, Tablet preserves lossless representations of real-world table images and spans a diverse set of tasks, including table-to-text generation (Parikh et al., 2020), table fact verification (Chen et al., 2020b), and table-based question answering (Pasupat & Liang, 2015). Tablet facilitates training and evaluation on table formats that align with those encountered in downstream applications, particularly in pixel-based agents that process visual input directly. It contains 4 million examples across 20 tasks, derived from 2 million unique tables, 88% of which retain their original visualization. To promote flexibility and foster future research, we provide both image and HTML representations of tables, together with metadata and links to the source datasets. Pairing table images with HTML enables models to learn from naturally occurring and synthetic table images when the former are not available. Tablet’s large-scale and task aggregation provide a unique, foundational resource that complements research integrating textual and visual modalities for training and reasoning (Liu et al., 2025; Jiang et al., 2025). We experimentally demonstrate that fine-tuning on Tablet improves performance on in- and out-of-domain benchmarks. Our work makes the following contributions:
  • Tablet addresses a critical gap in large-scale, visually faithful datasets for VTU, with 4 million examples and 2 million tables, most of which are preserved in their original form.
  • We show that supervised fine-tuning on Tablet  enhances VLM performance across seen and unseen table understanding tasks. We release a VLM model trained on Tablet that achieves state-of-the-art results on 10 out of 14 table understanding benchmarks.
  • Tablet  is designed for extensibility: all tasks are provided in a unified format, paired with HTML table serializations that can be rendered back into the original visualization, along with example metadata and links to the source datasets to facilitate reuse and reformulation.
Figure 1: Previous datasets render table images from serialized tables, losing original visual details. In contrast, Tablet locates and retrieves the original table visualizations across 14 tabular datasets, resulting in 4M examples grounded in 2M unique tables.
2 Related Work
Early work approached the problem of table understanding from a unimodal perspective, treating tables as structured text and developing general-purpose models that generalized across multiple tasks (Li et al., 2023; Zhang et al., 2024). Recent progress in natural image understanding afforded by increasingly better VLMs (Radford et al., 2021; Liu et al., 2023) has led to development of models that can successfully process visually rendered text beyond natural imagery and perform table understanding holistically, within a multimodal framework (Kim et al., 2022; Lee et al., 2023; Zhang et al., 2023; Ye et al., 2023; Hu et al., 2024; Alonso et al., 2024; Jiang et al., 2025; Su et al., 2024; Zhao et al., 2024; Zheng et al., 2024b; Su et al., 2024).
To this effect, several benchmarks have been proposed recently to evaluate progress in VTU. For instance, TableVQA-Bench (Kim et al., 2024) focuses on visual table question answering but relies on synthetic table images. In contrast, MMTBench (Titiya et al., 2025) includes original table visualizations, along with visually rich images and interleaved charts. While both are valuable for evaluation, they are limited in scope and do not provide large-scale training data, restricting their utility for developing generalizable models.
Larger training datasets (with training splits) have historically focused on image-based Table Structure Recognition (TSR), including PubTabNet (Zhong et al., 2020), TableBank (Li et al., 2020), and TabComp (Gautam et al., 2025). Beyond TSR, Alonso et al. (2024) created image-rendered versions of existing datasets like ToTTo (Parikh et al., 2020), relying on lossy renderings that discard visual features. WikiDT (Shi et al., 2024) preserves original table visualizations but only targets Visual TableQA. MMTab (Zheng et al., 2024b) is a large VTU dataset, with 150k TSR pre-training examples and 232k instruction examples across 19 tasks. While MMTab is a useful resource that we build upon, it relies on synthetic renderings of serialized tables, lacks traceability to original data sources, and remains limited in size compared to the scale needed for general-purpose VTU.
In this work, we do not introduce a task-specific dataset or model. Instead, we create a large-scale resource aimed at enhancing table understanding in general-purpose vision-language models like Gemma-3 (Team et al., 2025) or Qwen2-VL (Wang et al., 2024), under the assumption that tables for many tasks are seen as images, and therefore can be naturally processed by these models. This distinguishes Tablet from many recent efforts that introduce task-specific datasets or specialized architectures as it is designed to be a holist resource for advancing VLM capabilities for tables.
3 The Tablet Dataset
3.1 Overview
Tablet is a large-scale dataset that aggregates a total of 4,066,545 examples, combining 20 TU tasks sourced from 14 datasets (see Section 3.3 for a breakdown of examples per task). Tablet includes 2,031,256 unique table images with visualizations from Wikipedia (61.5%), PubTabNet (25.1%; Zhong et al. 2020), TabMWP (1.9%; Lu et al. 2023, and synthetically rendered images (11.5%) from serialized tables (see Section 3.2 for details).
Examples in Tablet are framed as instructions, leveraging the benefits of instruction-tuning for large language models.Tablet does not introduce new data instances; rather it repurposes examples from existing datasets, referred to as seed datasets, rephrasing their tasks into an instruction format. Examples are drawn from 14 seed datasets: TURL (Deng et al., 2020), ToTTo (Parikh et al., 2020), TabFact (Chen et al., 2020a), WikiTableQuestions (Pasupat & Liang, 2015), HybridQA (Chen et al., 2020b), HiTab (Cheng et al., 2022), PubTabNet (Zhong et al., 2020), TabMWP (Lu et al., 2023), TAT-QA (Zhu et al., 2021), InfoTabs (Gupta et al., 2020), WikiBIO (Lebret et al., 2016), FeTaQA (Nan et al., 2022), MMTab Zheng et al. (2024b), and DocStruct4M (Hu et al., 2024).
Datasets that include only images and fixed prompts restrict researchers to the provided visualizations and instruction formulations, limiting the exploration of new prompting techniques or the use of table data to create additional examples. To avoid this, each example in Tablet is provided in a unified format, including the instruction used in this work, its corresponding atomic data, the HTML version of the table, the original source ID, and additional metadata (detailed in Appendix7.4).
Task Seeds Train Dev Test
ent_link TURL 1,236,128 (35.3%) 74,282 (35.4%) 213,494 (60.8%)
col_type TURL 602,406 (17.2%) 13,188 (6.3%) 12,802 (3.6%)
struct_aware_parse M3 513,482 (14.6%) 9,115 (4.3%) 1,102 (0.3%)
wikibio WikiBIO 582,659 (16.6%) 72,831 (34.7%) 72,831 (20.7%)
hybridqa HybridQA 62,670 (1.8%) 3,466 (1.7%) 3,463 (1.0%)
fetaqa ToTTo 3,006 (0.1%) 577 (0.3%) 1,079 (0.3%)
hitab NSF, StatCan, ToTTo 7,417 (0.2%) 1,670 (0.8%) 1,584 (0.5%)
infotabs InfoTabs 16,538 (0.5%) 1,800 (0.9%) 5,400 (1.5%)
tabfact TabFact 87,717 (2.5%) 12,389 (5.9%) 12,326 (3.5%)
tabmwp TabMWP 23,059 (0.7%) 7,686 (3.7%) 7,686 (2.2%)
tat-qa TAT-QA 2,201 (0.1%) 278 (0.1%) 277 (0.1%)
totto ToTTo 110,934 (3.2%) 7,077 (3.4%) 7,084 (2.0%)
wikitq WikiTableQuestions 14,152 (0.4%) 3,537 (1.7%) 4,344 (1.2%)
rel_extraction TURL 60,615 (1.7%) 2,145 (1.0%) 2,030 (0.6%)
table_instruction M1 136,944 (3.9%) 0 (0.0%) 0 (0.0%)
row_column_extraction M1 7,721 (0.2%) 0 (0.0%) 957 (0.3%)
table_cell_extraction M1 7,727 (0.2%) 0 (0.0%) 966 (0.3%)
table_cell_location M1 7,708 (0.2%) 0 (0.0%) 956 (0.3%)
table_recognition M1 6,927 (0.2%) 0 (0.0%) 912 (0.3%)
table_size_detection M1 7,800 (0.2%) 0 (0.0%) 950 (0.3%)
merged_cell_detection M2 7,500 (0.2%) 0 (0.0%) 950 (0.3%)
Total 3,505,311 210,041 351,193
Table 1: Tablet splits, tasks, and seed datasets; M1: InfoTabs, NSF, StatCan, TabMWP, TAT-QA, TabFact, ToTTo, WikiBIO, WikiTableQuestions; M2: InfoTabs, NSF, StatCan, TAT-QA, TabFact, ToTTo, WikiBIO; M3: PubTabNet, TabFact, WikiTableQuestions. Proportion of examples in each set (train/dev/test) is shown in gray within parentheses.
3.2 Image sources
Wikipedia Tables Wikipedia tables generally follow a hierarchical table structure and are often rich in visual information. 61.5% of the tables in Tablet are lossless visualizations sourced from Wikipedia (See Appendix 7.2 for a breakdown of image sources per task). These tables are referenced in 67.4% of all examples in the dataset. Each table from the seed datasets was traced back to its original visualization in the corresponding Wikipedia article and captured in a screenshot. [^3] Table images were rendered using Firefox in headless mode (version 142.0.1 with GeckoDriver version 0.36.0) at an effective rendering density of 96 PPI (which stands for Pixels Per Inch).. A key challenge in retrieving the original visualizations was that the seed datasets were created at different times, while Wikipedia articles are continuously updated. To address this, we relied on the metadata released with the seed datasets and also reached out to their authors to find out when the tables were harvested (see Appendix 7.3). We used Wikipedia’s archiving API to retrieve each article as it appeared at the time of crawling. All Wikipedia tables in Tablet are linked to their source via both the page identifier and the corresponding revision identifier (oldid). As Wikipedia pages often contain multiple tables, we needed to identify which one corresponded to the seed example. We did this by computing the Levenshtein (1966) edit distance between Wikipedia tables and seed tables, both represented in markdown format, and selecting those with the highest similarity value. We set a minimum similarity threshold of 0.70; in cases without a match, the serialized seed table was rendered as an image.
Synthetic Tables Wikipedia tables that could not be retrieved, either due to low similarity scores or other issues, were synthetically generated by converting their serialized format to HTML and rendering them using the same browser configuration used for the original Wikipedia tables. These synthetic tables account for 11.5% of the dataset, contributing 747,002 additional examples (18.4% of the total). In Section 5, we show that combining original table images with synthetic visualizations during training allows for an increase in example count and visualization diversity that leads to better model performance compared to using synthetic or original tables alone.
Document Tables We also incorporate tables from other domains to improve model generalization. These include tables from scientific articles in PubTabNet (Zhong et al., 2020) and rendered tables from TabMWP (Lu et al., 2023). PubTabNet contributes 509,892 tables to our dataset (25.1% of the total), accounting for 97.5% of the structure-aware learning task incorporated from DocStruct4M (Hu et al., 2024). TabMWP adds 38,431 tables (1.9% of the total); while smaller in scale, TabMWP offers valuable coverage of numerical reasoning examples.
3.3 Table Understanding Tasks
Tablet aggregates a total of 20 TU tasks grouped into 7 broad categories: Table Interpretation, Question Answering, Table-to-Text Generation, Table Numerical Reasoning, Table Natural Language Inference (NLI), and Table Structure Understanding. Subtasks within each category require different skills and demonstrate various degrees of complexity, ranging from basic table structure understanding to downstream tasks that combine tabular reasoning with other skills, such as information retrieval, numerical reasoning, or natural language inference. Table 1 provides a breakdown of examples per task and their source dataset. We briefly describe each task below and provide more details in Appendix 7.5.
Table Interpretation comprises three tasks, namely Column Type Annotation, Entity Linking, and Relation Extraction (ent_link, col_type, and rel_extraction in Table 1). Sourced from TURL (Deng et al., 2020), these tasks do not represent downstream applications but target basic table understanding skills essential for tackling more complex tasks. All instructions for these tasks were aggregated from TURL. They constitute the largest group in our training set, with 1,899,149 examples (54.2%).
Table Question Answering is represented by Free-form Table QA (FeTaQA, Nan et al. 2022), Hierarchical Table QA (HiTabQA; Cheng et al. 2022), Hybrid QA (HybridQA; Chen et al. 2020b), and Table QA (WikiTableQuestions; Pasupat & Liang 2015). These tasks range from simple question answering over table content to multi-hop reasoning that combines textual and tabular information. They also vary in terms of table complexity and the expected answer format. Question answering represents 2.5% of Tablet’s training set (87,245 examples).
Table-to-Text Generation is exemplified by two subtasks, namely Highlighted Cell Text Generation (ToTTo; Parikh et al. 2020) and Wikipedia Biography Generation (WikiBio; Lebret et al. 2016). Both tasks involve generating text based on table content, concise biographies in the case of WikiBio, and short descriptions based on visually highlighted cells, in the case of ToTTo. Together, they account for 693,593 training examples (19.8%).
Table Numerical Reasoning combines Tabular Math Word Problem Solving (TabMWP; Lu et al. 2023) and Hybrid-context Financial QA (TAT-QA; Zhu et al. 2021). Given a table and a mathematical question, the model must perform reasoning over table values. This type of numerical reasoning represents 0.8% of Tablet’s training set (25,260 examples).
Table NLI includes two entailment tasks: Table Fact Verification (TabFact; Chen et al. 2020a) and Infobox Natural Language Inference (InfoTabs; Gupta et al. 2020) where statements as supported or refuted based on table content. They represent 104,255 examples (3%) in Tablet’s training set.
Table Structure Understanding is an umbrella category for tasks designed to facilitate structural understanding of tables: Merged Cell Detection, Row and Column Extraction, Table Cell Extraction, Table Cell Location, Table Recognition, Table Size Detection, and Structure-aware Parsing in Table 1). Examples for the first seven tasks are aggregated from MMTab, while Structure-aware Parsing is derived from DocStruct4M (Hu et al., 2024). These tasks are based on tables from multiple seed datasets (see rows with seeds M2, M1, and M3), including InfoTabs (Gupta et al., 2020), NSF (National Center for Science and Engineering Statistics, 2019), StatCan (Statistics Canada, 2024), TabMWP (Lu et al., 2023), TAT-QA (Zhu et al., 2021), TabFact (Chen et al., 2020a), ToTTo (Parikh et al., 2020), WikiBio (Lebret et al., 2016), WikiTableQuestions (Pasupat & Liang, 2015), and PubTabNet Zhong et al. (2020). While we maintain the original MMTab instructions, we use our own table visualizations. This group contributes 558,865 examples (15.8%) to our training.
Instruction Following To facilitate evaluation, we fine-tune and evaluate all models using instructions that explicitly require the final answer to be encapsulated in a JSON object, regardless of any additional tokens the model may generate. To increase instruction diversity and mitigate catastrophic forgetting, we include a dedicated instruction-following set from MMTab. These instructions rephrase a subset of examples from the above tasks using a different template and do not require the JSON output format. While not a task in itself, this set adds instruction diversity and reduces overfitting to a single output style. We aggregate these examples directly from MMTab while using our visualizations. This set contributes 136,944 training examples (3.9%).
3.4 Linking to Sources and Highlighting
Tablet is designed with extensibility in mind. Existing datasets such as TableInstruct (Zhang et al., 2024), DocStruct4M (Hu et al., 2024), and MMTab (Zheng et al., 2024b) do not provide a clear reference to the original examples from which their instructions were derived. However, the absence of such pointers, makes reuse, modification, and augmentation difficult. Additionally, image-based datasets such as MMTab lack serialized table versions in text format, providing only rendered images as representations. To address these issues, each example in Tablet includes both the identifier of the original example and the identifier of the corresponding table from the source dataset. Identifiers follow a simple, human-readable format, and the released code provides a function to retrieve the original examples given their IDs.
To support future extensions and new tasks, Tablet also provides serialized versions of its tables in HTML. For the tables successfully retrieved from Wikipedia, this HTML corresponds to the raw table object in the original HTML of the article. These tables can be rendered with their full visualization using Wikipedia’s official CSS stylesheets or through the rendering functions provided with our code. The remaining tables are either provided as HTML in the seed datasets or converted into HTML from their serialized representations.
This feature is particularly useful in tasks involving cell pointing or highlighting. Prior work has shown that models can achieve competitive results by relying uniquely on highlighted values (An et al., 2022; Alonso et al., 2024). We also observed in our experiments that as long as highlighted cell values are mentioned in the prompt, models are able to perform the task (e.g., Column Type Annotation, Entity Linking, Relation Extraction, FeTaQA, and ToTTo) regardless of the table provided. In other words, the model learns to ignore the table during training, which can be detrimental for tasks requiring genuine table understanding. We mitigate this, by exploiting the HTML serialization to directly locate the highlighted cells in the source tables. By cross-referencing the highlighted spans from the seed datasets with the actual table content, we generate versions of the tables with explicit highlights while preserving the original visual features.
3.5 Comparison with MMTab
MMTab (Zheng et al., 2024b) is also a dataset for visual table understanding; however, unlike Tablet, it does not include the full set of training examples from the original datasets (e.g., only 12.4% of ToTTo compared to our 91.8%; see Appendix 7.6), and it provides no development splits. MMTab covers 20 tasks while Tablet covers 21 (inculding instruction tuning); four tasks are unique to the former and five to the latter, and overall our dataset is 9.36 times larger (comprising 4,066,545 examples compared to MMTab’s approximate 433,376 examples). Moreover, only a few tasks in MMTab include identifiers that can be linked back to their source examples, limiting developers to the provided instruction text. While MMTab relies on synthetically rendered tables with predefined styles, which omit original formatting and embedded images, Tablet retains the original visualizations, enabling models to exploit a fuller range of visual features.
4 Experimental Setting
Our experiments are motivated by two questions: (1) Does training VTU models with lossless table visualizations improve performance on VTU tasks? and (2) Is Tablet diverse enough so that supervised fine-tuning (SFT) improves generalization on unseen VTU tasks? We also examine whether training only with original visualizations performs better than mixing original and synthetic ones, whether a balanced task distribution is preferable to the full dataset distribution, and whether including various table interpretation tasks contributes to VTU performance.
Backbone Model All experiments employ the same backbone model, namely Qwen2.5-VL with 7B parameters (Qwen et al., 2025). This model provides a good tradeoff between performance and computational requirements. Together with InternVL3 (Zhu et al., 2025), it was the best-performing open-weight VLM in its class, setting a new benchmark for Single Page Document VQA (Mathew et al., 2021) at the time of writing. Exploring the best backbone model for VTU or advanced SFT strategies is outside the scope of this work. Full SFT configurations are detailed in Appendix 8; results for other open-weight models on Tablet (test set) are reported in Appendix 8.4.
Evaluation We evaluate fine-tuned models on eight held-in tasks: three focus on Table QA (WikiTableQuestions, HiTabQA, FeTaQA), two on text generation (WikiBio, ToTTo), two on Table Numerical Reasoning tasks (TabMWP, TAT-QA), and one on Table NLI (TabFact). To test generalization, we further evaluate on six held-out tasks: HybridQA, InfoTabs, TabMCQ (Jauhar et al., 2016), AIT-QA (Katsis et al., 2022), Table Recognition, and PubHealthTab (Akhtar et al., 2022). The last four are directly extracted from MMTab and use instructions generated with their own templates, adding diversity in both table and instruction styles. Although included in Tablet, the 86,135 training examples from HybridQA, InfoTabs, and Table Recognition were excluded from fine-tuning to enable held-out evaluation. Models trained on the full dataset would potentially achieve even stronger results than those reported here.
5 Results and Analysis
[…]
WikiBio ToTTo WikiTQ TabMWP HiTAB TabFact FeTaQA TAT-QA
Model synth org synth org synth org org synth org synth org synth org org
0-Shot 2.0 2.5 10.8 8.9 57.7 50.8 51.9 11.1 2.5 70.4 59.4 9.9 9.8 48.7
[…]
Table 2: Comparison of Qwen2.5-VL-7B-Instruct models fine-tuned on different Tablet-B variants and in zero-shot mode; models are evaluated on original (lossless) table visualizations and synthetic ones across held-in benchmarks. Results for TabMWP and TAT-QA exclude the mixed setting, as all tables already use original visualizations. We report exact match accuracy for WikiTQ, TabMWP, TabFact, and TAT-QA; BLEU for WikiBio, ToTTo, and FeTaQA; and F1 for HiTab. ToTTo results are reported on its development set. Best model per setting (synth/org) shown in bold.
Our results in Table 2 (held-in tasks) show that lossless models outperform synthetic ones in two tasks (WikiBio, HiTab), match performance in four (WikiTQ, TabMWP, FeTaQA, TAT-QA), and underperform in one (TabFact). For reference, we also report the 0-shot performance of Qwen2.5-VL-7B. The fact that there is no pronounced difference between lossless and synthetic visualizations is perhaps not surprising since most benchmarks were originally designed around serialized, text-only representations, without paying heed to visual cues (TURL, WikiBIO, ToTTo, TAT-QA). Thus, while lossless visualizations contain more information, current benchmarks do not test whether models benefit from them. Developing benchmarks with high visual variability and challenging layouts would better highlight scenarios where lossless visualizations offer a distinct advantage.
[…]
Performance on Unseen VTU Tasks with Tablet We next evaluate whether Tablet’s task diversity improves performance on unseen datasets. While fine-tuned models are expected to perform better on tasks seen during training, improved performance on held-out datasets would suggest that training on Tablet enhances the model’s VTU capabilities. Tables 3 and 4 report our results on held-in and held-out tasks, respectively. We present comparisons between zero- and one-shot Qwen2.5-VL-7B [^4] We should note that Qwen2.5-VL-7B is a competitive baseline that has undergone extensive pre-training and instruction-tuning, achieving top scores in many TU and Visual Document Understanding benchmarks. and its fine-tuned instantiations trained on MMTab and different sizes of Tablet. For MMTab, all Wikipedia-sourced images are synthetic (61.8% of all tables in their dataset), whereas all Tablet versions use a mix of original (88%) and synthetic (12%) (See Appendix 7.2 for the distribution of image sources per task). For now, we focus solely on Tablet-L (a shorthand for Tablet-Large), which is our biggest dataset comprising 4M examples (we discuss smaller sizes in the next sections).
Held-in Datasets
WikiBio ToTTo WikiTQ TabMWP Hitab TabFact FeTaQA TAT-QA
0-Shot 2.5  9.1  53.4  59.1  31.2  73.9   7.0   6.9
1-Shot 4.4  16.1  49.8  57.7  37.3  72.8   8.5  10.1
MMTab 6.4  12.6  48.7  80.4  41.5  57.0   1.7   9.4
[…]
Table 3: Evaluation on eight held-in datasets  with Qwen2.5-VL-7B-Instruct in 0/1-shot mode and fine-tuned on MMTab and different Tablet sizes. Results for ToTTo correspond to evaluation on the development set. We report exact match accuracy for WikiTQ, TabMWP, TabFact, and TAT-QA; BLEU for WikiBio, ToTTo, and FeTaQA; and F1 for HiTab. Best performing model per task is shown in bold; we mark with $\ast$ models significantly different ($p<0.05$, using bootstrap resampling) from those trained on Tablet (highlighted in gray). Models that perform significantly better within the Tablet group are underlined.
Held-out-Datasets
Infotabs TabMCQ AIT-QA PubHealthTab HybridQA TabRec
0-Shot  64.5 84.4  51.7  64.7 34.2  24.5
1-Shot 65.5  61.0  43.1  70.0  30.5  27.9
MMTab  56.9 89.1  56.6  63.9 27.1  43.6
[…]
Table 4: Left: Evaluation on five held-out datasets with Qwen2.5-VL-7B-Instruct (0/1-shot) and fine-tuned on MMTab and different Tablet sizes. We report Tree-Edit- Distance-based Similarity for Table Recognition (TabRec) and accuracy for all other benchmarks. HybridQA results are reported on the development set. Best performing model per task is shown in bold; we mark with $\ast$ models significantly different ($p<0.05$, using bootstrap resampling) from those trained on Tablet (highlighted in gray). Models that perform significantly better within the Tablet group are underlined. Right: Distribution of examples per task in Tablet-L (b) and Tablet-M (a).
[Uncaptioned image]       a.Tablet-M         b. Tablet-L
We observe that the model fine-tuned on Tablet-L  performs overwhelmingly better on held-in and held-out datasets. It dominates in all QA benchmarks except HybridQA, where long-context proficiency is likely reduced since its training set was excluded from fine-tuning. Notably, in tasks such as HiTab, FeTaQA, and TAT-QA, where Tablet includes fewer training examples than MMTab, the model still performs considerably better, suggesting benefits from transfer across tasks, original visualizations, or simply due to a higher-quality subset of examples. Finally, tasks that include hierarchical table structures such as ToTTo and HiTab show clear gains. Table cell hierarchy is not lost in synthetic tables, but it is often conveyed more clearly in their original visualizations. Our results are encouraging, particularly when considering that portions of the held-out datasets, including test sets, may have been encountered during petraining. Models trained on Tablet outperform 0/1-shot Qwen2.5-VL-7B on 12 out 14 (held-in and held-out) tasks.
Interestingly, our gains extend beyond unseen tasks. For instance, ToTTo, HiTab, TabFact, FeTaQA, and TAT-QA contain the same number of training examples across all Tablet variants, yet performance consistently improves as additional tasks are incorporated. This suggests that fine-tuning on a broader task set facilitates transfer learning, yielding improvements even on tasks where the model was already competitive.
Optimal Tablet Size Since training on 4M examples is resource-intensive, and tasks in Tablet-L exhibit substantial variation in size (see (b) pie chart in Table 4), we evaluate whether a smaller but more balanced dataset might perform comparably. We create Tablet-M (a shorthand for Tablet-Medium) by capping each task at 140k examples. For benchmarks exceeding this cap (Column Type Annotation, Entity Linking, WikiBio), a random 140k subset is sampled and included as representative for that task. This results in a dataset with 1,117,217 training examples. We exclude HybridQA, InfoTabs, and Table Recognition from the training set to allow for held-out evaluation, leaving 1,031,082 examples in Tablet-M and 3,419,176 in the full dataset of our experiments. The capped distribution is shown in pie chart (b), Table 4 and further details on training sets are in Appendix 7.1.
The model trained on Tablet-M performs competitively and in some cases, even better than Tablet-L (see Table 4, rows in gray), offering a compelling trade-off between dataset size and model effectiveness. That said, the full dataset still yields slightly better performance overall and remains a valuable resource for future work that can benefit from large-scale, lossless training.
The Benefit of Training on Table Interpretation Tasks Table Interpretation tasks represent 54.2% of the training examples in Tablet-L, making them the largest category by volume and the second longest in instruction length, following HybridQA (see Appendix 8.1). While Deng et al. (2020) demonstrated the benefits of these tasks for textual TU, it is unclear whether this carries over to VTU. If these tasks prove unhelpful, removing them would save significant resources.
We therefore remove all Table Interpretation tasks from Tablet-M and fine-tune on this smaller version (we use Tablet-S as a shorthand for Tablet-Small) which excludes any tasks related to Column Type Annotation, Entity Linking, and Relation Extraction resulting in a dataset with a total of 690,467 training examples (67% of the example count in Tablet-M). As shown in Table 4, there are benefits to be gained from including these tasks. Models trained on Tablet-M outperform those trained on Tablet-S on six out of 8 held-in benchmarks (Table 3) and achieve comparable performance on the remaining two. Similar gains are observed in four out of five held-out tasks (Table 4). These findings demonstrate that the benefits observed in the textual domain (Deng et al., 2020) extend to multimodal VTU.
6 Conclusions
In this work, we introduced Tablet, a large-scale dataset for Visual Table Understanding that aggregates 4 million examples across 20 tasks and 2 million unique tables. Unlike prior resources, Tablet preserves original table visualizations whenever possible, provides both HTML and image representations, and maintains full traceability to the source datasets. Extensive experiments showed that (1) training on Tablet increases robustness to real-world tables compared to synthetic renderings; (2) models fine-tuned on Tablet consistently outperform those trained on related VTU datasets across held-in and held-out tasks; and (3) Tablet brings improvements on unseen benchmarks, which suggests that its diversity supports transfer across VTU tasks. We hope Tablet will foster further research on Visual Table Understanding, including the design of new tasks and the evaluation of models in settings that reflect real-world tabular data.